OpenStudy (anonymous):

Solve the equation for x by first making an appropriate substitution, such as u=e^x 5e^2x-16e^x+3=0

OpenStudy (campbell_st):

let\[ u = e^x \] so \[e^{2x} = (e^x)^2\] using the power of a power index law so you will have \[5u^2 - 16u + 3 = 0\] which can be factorised.... (5u - 1)(u - 3) = 0 thats the hard part.. all you need to do is solve... \[e^x = \frac{1}{5}..... and ..... e^x = 3\] use logarithms...

OpenStudy (anonymous):

It says its not right for both of them =/

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

thats correct..its not right...as you still need to find x.... that is what value of x when a power of e is equal to 1/5 and 3 so you need to take the base e log of both sides to get your answer... one value will be positive and one value will be negative

OpenStudy (anonymous):

Ahhh! I got it finally...thank you!